OpenStudy (valirium):

What is the vertex form of the equation? y = -x^2 + 12x - 4

Directrix (directrix):

The vertex form of a quadratic is of the format: y = a*(x-h)^2 + k where (h,k) is the vertex of the parabola. "a" is a constant which we have to figure out to solve this problem.

OpenStudy (valirium):

So I have to graph something for this problem?

OpenStudy (valirium):

Sorry, I'm not good at these kinds of problems.

Directrix (directrix):

I think we're going to have to complete the square technique on the given equation to get the vertex form.

Directrix (directrix):

y = -x^2 + 12x - 4 y = - ( x^2 - 12x + __?__) - 4 + __?__

Directrix (directrix):

@Valirium What is one-half of negative 12 ?

OpenStudy (valirium):

6?

OpenStudy (valirium):

ummm -4 and 3?

OpenStudy (valirium):

Idk I'm confused

Directrix (directrix):

Okay, we're on the way. What is one-half of negative 12 ? That is negative 6.

Directrix (directrix):

Any questions?

OpenStudy (valirium):

I got it. Thank you for helping! :)

Directrix (directrix):

You are welcome.